# Black Hole Cold Brew: Fermi Degeneracy Pressure

### v1

@@ Abstract
We investigate the dynamical instability of a self-gravitating thermal system in the quantum regime, where Fermi degeneracy pressure becomes significant. Using a truncated Fermi–Dirac distribution and solving the Tolman–Oppenheimer–Volkoff equation, we identify marginally stable configurations following Chandrasekhar’s criterion. While Fermi pressure stabilizes a system against gravitational collapse in Newtonian gravity, in general relativity it can instead drive the instability, enabling collapse even at low temperatures. We discuss implications for the formation of massive black holes in the early Universe through the gravothermal collapse of dark matter.
@@ I Introduction
Black holes are an inevitable consequence of gravitational collapse as predicted by Einstein’s general relativity, and their existence have been confirmed in various astrophysical observations, including those associated with quasars Mortlock:2011wz ; Wu:2015vd ; Banados:2017unc ; Wang:2018tr ; Onoue:2019ty ; Matsuoka:2019vy ; Yang:2020ut ; Wang:2021ve ; Larson:2023aa , stellar dynamics Schodel:2002py ; Ghez:2008ms ; GRAVITY:2018ofz , gravitational-wave detections, and JWST Matthee:2023utn ; Harikane:2023aa ; Kocevski:2023aa ; Pacucci:2023oci ; Maiolino:2023bpi . In theory, stellar-mass black holes, such as those detected by LIGO via merger events, typically form from the death of massive stars LIGOScientific:2016aoc ; Olsen:2022pin ; KAGRA:2021duu ; LIGOScientific:2025rid . The origin of supermassive black holes with masses larger than $10^{6}{\rm\,M}_{\odot}$, particularly those in the early Universe, is more puzzling (see Inayoshi:2019fun ; Woods:2018lty ; Haemmerle:2020iqg ; Volonteri:2021sfo for reviews and references therein). A popular mechanism is that direct collapse of pristine gas in proto-galaxies into a massive seed black hole, without first undergoing star formation Rees:1984si ; Bromm:2002hb ; Begelman:2006db , and the seed further grows by accretion. Primordial black holes Carr:2025kdk ; Bird:2016dcv ; Sasaki:2016jop ; Jedamzik:2020omx ; Carr:2020gox ; Musco:2023dak ; Andres-Carcasona:2024wqk ; Feng:2024obn could also play a role in seeding supermassive black holes in the early Universe Davoudiasl:2021ijv ; Flores:2023zpf ; Lu:2024xnb ; Qin:2025ymc ; Zhang:2025tgm . Another intriguing mechanism is that if dark matter particles carry self-interactions Tulin:2017ara , the central regions of dark matter halos can undergo gravothermal collapse and form a seed black hole, such interactions could be elastic Balberg:2001qg ; Pollack:2014rja ; Feng:2020kxv ; Meshveliani:2022rih ; Jiang:2025jtr ; Roberts:2025poo ; Feng:2025rzf or inelastic Choquette:2018lvq ; Xiao:2021ftk ; Fernandez:2022zmc ; Shen:2025evo .
@@
It is important to investigate the dynamical conditions under which a seed black hole can form in light of these seeding mechanisms. The dynamical instability of self-gravitating systems within the framework of general relativity was first studied in the context of supermassive stars Chandrasekhar:1964zza ; Chandrasekhar:1964zzb ; Fowler:1964zza ; Bardeen:1965 ; Zeldovich:1966 ; Ipser:1980 . In general, a star tends to collapse into a black hole once its gravitational energy becomes comparable to its rest-mass energy, i.e., $GM^{2}/R\sim Mc^{2}$, where $G$ is Newton’s constant, $c$ is the speed of light, and $R$ is the characteristic radius of the star. In contrast to Newtonian gravity, where pressure acts as a “force” opposing collapse, in general relativity pressure is part of the energy–momentum tensor and therefore sources gravity, which in turn destabilizes the star. Additional factors, such as the equation of state, the cosmological constant and spatial dimensionality Boehmer:2005kk ; Bordbar:2015wva ; Posada:2020svn ; Feng:2021acd ; Feng:2022fuk , also play a role in determining stability.
@@ II The Fermi–Dirac Model
$f(\epsilon)=\left\{\begin{matrix}\frac{1-e^{(\epsilon-\epsilon_{c})/k_{B}T}}{e^{(\epsilon-\mu)/k_{B}T}+1},&\epsilon\leq\epsilon_{c};\\ 0,&\text{otherwise,}\end{matrix}\right$. (1)
@@
For convenience, we introduce the dimensionless quantities Feng:2021rst :
$x(r)$ $\equiv$ $\epsilon(r)/{k_{B}T(r)}$ (3)
$w(r)$ $\equiv$ ${\epsilon_{c}(r)}/{k_{B}T(r)}$ (4)
$\alpha(r)$ $\equiv$ ${\mu(r)}/{k_{B}T(r)}$ (5)
$b$ $\equiv$ ${k_{B}T(R)}/{mc^{2}}$ (6)
@@
In the limit $r=R$, we see that $w(R)=0$, i.e., the cutoff energy vanishes at the outer boundary as expected for a self-gravitating system. Consider $\alpha(R)=0$, such that particles in the sphere are completely degenerate $\alpha(r)>0$, the cutoff energy and the degeneracy take the same value as indicated in Eq. 7, $\epsilon_{c}(r)=\mu(r)>0$ is the energy required to move a particle from the outer boundary $R$ to the radius $r$. For the model we consider, we take a specific form of the truncation as in Eq. 1, but the relations shown in Eqs. 2 and 7 are generic as they are based on laws of thermodynamics and general relativity.
We can obtain the the profile of $w(r)$ by solving the TOV equation Feng:2021rst
@@
respectively, where $g=2s+1$ is the spin multiplicity of the gas particle, $h$ is the Planck constant. In addition, the characteristic size of the system can be determined by the fiducial length in numerical implementation,
@@
where $m_{\text{Pl}}=(\hbar c/G)^{1/2}$ is the Planck mass, $\lambda_{C}=\hbar/mc$ is the reduced Compton wavelength of the particle ($\hbar=h/2\pi$), and $\hat{r}$ is the dimensionless radius. Furthermore, we can express the dimensional quantities of the system using their corresponding dimensionless counterpart denoted with a “hat,” i.e., $n=(c^{2}/Gm\zeta^{2})\hat{n}$, $\rho=(c^{2}/G\zeta^{2})\hat{\rho}$, $u=(c^{4}/G\zeta^{2})\hat{u}$, $p=(c^{4}/G\zeta^{2})\hat{p}$ and $M=(c^{2}\zeta/G)\hat{M}$.
Figure 1: The normalized classical Maxwell–Boltzmann distribution (top) and quantum Fermi–Dirac distribution with $\alpha(R)=0$ (bottom). The cutoff energy is fixed at $bw=0.3$, and the boundary temperature varies as $b=k_{B}T(R)/mc^{2}=10^{-3}$ (blue), $10^{-2}$ (orange), and $10^{-1}$ (red). In the classical limit, the energy peak shifts toward higher energies as the temperature increases, whereas in the quantum limit, the trend reverses as the fermions become fully degenerate.
@@ III Fermi Degeneracy Pressure
In Fig. 1, we compare the normalized distribution function in the classical (top) and quantum (bottom) limits in the relativistic regime $bw=0.3$ for the boundary temperature $b=10^{-3}$ (blue), $10^{-2}$ (orange), and $10^{-1}$ (red). In the classical Maxwell-Boltzmann limit, most of the particles are non-relativistic with $bx\lesssim 10^{-2}$ for $b=10^{-3}$. As $b$ increases to $0.1$, more and more particles occupy the high energy state. Thus to trigger the instability in the this case, one must increase the boundary temperature $b$ such that the majority of the particles can in the high energy state $bx\sim 0.1$. On the other hand, for the quantum limit $\alpha(R)=0$, the distribution is peaked at $bx\approx 0.3$ even at the low temperature $b=10^{-3}$. As $b$ increases further, the distribution changes mildly as in this limit the Pauli blocking effect dominates.
@@
Furthermore, we evaluate the velocity dispersion $v/c=\sqrt{3p/\rho c^{2}}$ and $\gamma=1+p/u$, where pressure $p$, energy density $\rho c^{2}$, and internal energy $u$ are given in Eq. 9. We consider the Maxwell–Boltzmann distribution and the Fermi–Dirac distribution with $\alpha(R)=\mu(R)/k_{B}T(R)=-10$ and 0, ranging from the classical to quantum limits. The $b$ values are $b=k_{B}T(R)/mc^{2}=10^{-3}$, $10^{-2}$, and $10^{-3}$, as the same as those in Fig. 1.
@@ IV Dynamical Instability
In Fig. 3, we show the stable and unstable regions in the plane of the boundary temperature $b$ and the central cutoff $bw(0)$ for the degeneracy $\alpha(R)=-50$, 10-10, 5-5, and 0, covering the classical to quantum limits. In this classical limit ($\alpha(R)=-50$), the instability can only occur when the boundary temperature is high $b=k_{B}T(R)/mc^{2}>9\times 10^{-2}$ and the central cutoff is $bw(0)\gtrsim 0.35$ (relativistic regime, see Eq. 11). As the degeneracy $\alpha(R)$ increases to 10-10, another unstable region appears for $b<1.5\times 10^{-2}$ due to Fermi pressure, while the requirement on the central cutoff remains almost the same. In this case, the unstable region with high boundary temperature is still preserved, where the thermal pressure is the dominant source of the instability. Interestingly, there is a gap in the boundary temperature between the two unstable regions. In the gap region, both thermal pressure and Fermi pressure are relevant, but overall they are not strong enough to trigger the instability. As $\alpha(R)$ further increases, the gap shrinks and vanishes for $\alpha(R)\gtrsim-5$. For $\alpha(R)=0$, both regions join smoothly and the instability condition becomes insensitive to the boundary temperature $b$, as long as the central cutoff $bw(0)\gtrsim 0.3$.
@@
$\lambda_{\rm dB}(r)$ $=\lambda_{C}\left[\left(1+\frac{3k_{B}T(r)}{2mc^{2}}\right)^{2}-1\right]^{-1/2}$ (14)
$=\lambda_{C}\left[\left(1+\frac{3b/2}{1-bw(r)}\right)^{2}-1\right]^{-1/2},$
where $\lambda_{C}=\hbar/mc$ is the reduced Compton wavelength. Using Eq. 10, we have the following relation
@@ V Black hole mass
With our numerical results as tabulated in Appendix C, we can calculate the critical mass for marginally stable configurations satisfying $\left<\gamma\right>=\gamma_{\rm cr}$. From Eq. 10, the gravitational mass of the bound sphere is given by
@@
When the degeneracy is large, Fermi pressure is the dominant source to trigger the instability even a relatively low boundary temperature. In this regime, the critical mass is independent of the boundary temperature. In contrast, for the classical Maxwell–Boltzmann distribution, where thermal pressure is the only source of the instability and the only way to trigger the instability is to increase the boundary temperature. Note for $\alpha(R)=-50$, the system is in the classical limit and follows the Maxwell–Boltzmann distribution, there are no marginally stable configurations in the low-temperature regime as indicated in Fig. 3.
Figure 4: Critical mass $M$ as a function of particle mass $m$ for the degeneracy $\alpha(R)=-50$ (left), 10-10 (middle), and 0 (right), obtained by varying the boundary temperature $b=0.5$ (red), $10^{-1}$ (orange), and $10^{-2}$ (blue). At the low temperature $b=10^{-2}$, marginally stable configurations are not found for $\alpha(R)=-50$, where the classical Maxwell–Boltzmann distribution is recovered, while the critical mass converges for $\alpha(R)=-10$ and 0 as the core becomes completely degenerate.
Fig. 4 shows the critical mass as a function of particle mass for $\alpha(R)=-50$ (left), 10-10 (middle), and 0 (right) for $b=0.5$ (red), $0.1$ (orange), and $10^{-2}$ (blue). For a given $\alpha(R)$, the critical mass increases with the boundary temperature $b$. For $\alpha(R)=-50$, the instability cannot be triggered at a low temperature $b=10^{-2}$, but this case has the highest critical mass at the high-temperature regime among the three examples. Thus in the classical regime, black hole seeds tend to more massive. In the low-temperature regime $b=10^{-2}$, the critical mass converges and becomes insensitive to $\alpha(R)$. In this completely degenerate regime, the crtical mass reach the minimum for a given particle mass.
@@
Now we consider scenarios where dark matter clumps collapse in to black holes due to gravothermal evolution, such as in the self-interacting dark matter framework Balberg:2001qg ; Feng:2020kxv ; Feng:2021rst . For fermonic dark matter, there is a lower bound on the particle mass, i.e., the Tremaine-Gunn bound Tremaine:1979we , below which dark matter halos cannot form. Recent studies found that the lower bound is $m\gtrsim 0.4\,{\rm~keV}/c^{2}$ Boyarsky:2008ju , although it could be relaxed if many distinct species of fermions are allowed Davoudiasl:2020uig . Taking $m=0.4\,{\rm keV/c^{2}}$ in Eq. 18, we find that the critical mass is $3.9\times 10^{12}\,{\rm M_{\odot}}$. This would be larger than the most massive supermassive black hole observed so far. For a thermally produced dark matter particle, the gravitational lensing constraint from JWST data set a lower limit on its mass of $6.1\,{\rm keV}/c^{2}$ Keeley:2024brx , corresponding to a critical mass of $1.67\times 10^{10}\,{\rm M_{\odot}}$ from Eq. 18. Conversely, if $\sim 10^{7}\,{\rm M_{\odot}}$ black holes associated with the JWST little red dots Matthee:2023utn ; Harikane:2023aa ; Kocevski:2023aa ; Pacucci:2023oci ; Maiolino:2023bpi are formed through the collapse of degenerate fermionic cores, the particle mass is estimated as $m\sim 100{\rm\,keV/c^{2}}$ Arguelles:2023hab ; Arguelles:2023kqw .

### v2

@@ Abstract
We investigate the dynamical instability of a self-gravitating thermal system in the quantum regime, where Fermi degeneracy pressure becomes significant. Using a truncated Fermi–Dirac distribution and solving the Tolman–Oppenheimer–Volkoff equation, we identify marginally stable configurations following Chandrasekhar’s criterion. While Fermi pressure stabilizes a system against gravitational collapse in Newtonian gravity, in general relativity it can instead drive the instability, enabling collapse even at low temperatures. In the low-temperature limit, the critical mass is independent of the boundary temperature. We discuss implications for the formation of massive black holes in the early Universe through the gravothermal collapse of dark matter.
@@ I Introduction
Black holes are an inevitable consequence of gravitational collapse as predicted by Einstein’s general relativity, and their existence have been confirmed in various astrophysical observations, including those associated with quasars Mortlock:2011wz ; Wu:2015vd ; Banados:2017unc ; Wang:2018tr ; Onoue:2019ty ; Matsuoka:2019vy ; Yang:2020ut ; Wang:2021ve ; Larson:2023aa , stellar dynamics Schodel:2002py ; Ghez:2008ms ; GRAVITY:2018ofz , gravitational-wave detections, and the James Webb Space Telescope (JWST) Matthee:2023utn ; Harikane:2023aa ; Kocevski:2023aa ; Pacucci:2023oci ; Maiolino:2023bpi . In theory, stellar-mass black holes, such as those detected by LIGO via merger events, typically form from the death of massive stars LIGOScientific:2016aoc ; Olsen:2022pin ; KAGRA:2021duu ; LIGOScientific:2025rid . The origin of supermassive black holes with masses larger than $10^{6}{\rm\,M}_{\odot}$, particularly those in the early Universe, is more puzzling (see Inayoshi:2019fun ; Woods:2018lty ; Haemmerle:2020iqg ; Volonteri:2021sfo for reviews and references therein). A popular mechanism is the direct collapse of pristine gas in proto-galaxies into a massive seed black hole, without first undergoing star formation Rees:1984si ; Bromm:2002hb ; Begelman:2006db , and the seed further grows by accretion. Primordial black holes Carr:2025kdk ; Bird:2016dcv ; Sasaki:2016jop ; Jedamzik:2020omx ; Carr:2020gox ; Musco:2023dak ; Andres-Carcasona:2024wqk ; Feng:2024obn could also play a role in seeding supermassive black holes in the early Universe Davoudiasl:2021ijv ; Flores:2023zpf ; Lu:2024xnb ; Qin:2025ymc ; Zhang:2025tgm . Another intriguing mechanism is that if dark matter particles carry self-interactions Tulin:2017ara , the central regions of dark matter halos can undergo gravothermal collapse and form a seed black hole, such interactions could be elastic Balberg:2001qg ; Pollack:2014rja ; Feng:2020kxv ; Meshveliani:2022rih ; Jiang:2025jtr ; Roberts:2025poo ; Feng:2025rzf or inelastic Choquette:2018lvq ; Xiao:2021ftk ; Fernandez:2022zmc ; Shen:2025evo .
@@
It is important to investigate the dynamical conditions under which a seed black hole can form in light of these seeding mechanisms. The dynamical instability of self-gravitating systems within the framework of general relativity was first studied in the context of supermassive stars Chandrasekhar:1964zza ; Chandrasekhar:1964zzb ; Fowler:1964zza ; Bardeen:1965 ; Zeldovich:1966 ; Ipser:1980 . In general, a star tends to collapse into a black hole once its gravitational energy becomes comparable to its rest-mass energy, i.e., $GM^{2}/R\sim Mc^{2}$, where $G$ is Newton’s constant, $c$ is the speed of light, and $R$ is the characteristic radius of the star. In contrast to Newtonian gravity, where pressure acts as a “force” opposing collapse, in general relativity pressure is part of the energy-momentum tensor and therefore sources gravity, which in turn destabilizes the star. Additional factors, such as the equation of state, the cosmological constant and spatial dimensionality Boehmer:2005kk ; Bordbar:2015wva ; Posada:2020svn ; Feng:2021acd ; Feng:2022fuk , also play a role in determining stability.
@@ II The Fermi–Dirac Model
$f(\epsilon)=\begin{cases}\displaystyle\frac{1-e^{(\epsilon-\epsilon_{c})/k_{B}T}}{e^{(\epsilon-\mu)/k_{B}T}+1},&\epsilon\leq\epsilon_{c},\\[4.0pt] 0,&\text{otherwise.}\end{cases}$ (1)
@@
For convenience, we introduce the dimensionless quantities Feng:2021rst ,
$x(r)$ $\equiv$ $\epsilon(r)/{k_{B}T(r)},$ (3)
$w(r)$ $\equiv$ ${\epsilon_{c}(r)}/{k_{B}T(r)},$ (4)
$\alpha(r)$ $\equiv$ ${\mu(r)}/{k_{B}T(r)},$ (5)
$b$ $\equiv$ ${k_{B}T(R)}/{mc^{2}}.$ (6)
@@
In the limit $r=R$, we see that $w(R)=0$, i.e., the cutoff energy vanishes at the outer boundary as expected for a self-gravitating system. Consider $\alpha(R)=0$, such that particles in the sphere are completely degenerate $\alpha(r)>0$, the cutoff energy and the degeneracy take the same value as indicated in Eq. 7, and $\epsilon_{c}(r)=\mu(r)>0$ is the energy required to move a particle from the outer boundary $R$ to the radius $r$. For the model we consider, we take a specific form of the truncation as in Eq. 1, but the relations shown in Eqs. 2 and 7 are generic as they are based on laws of thermodynamics and general relativity.
We can obtain the profile of $w(r)$ by solving the TOV equation Feng:2021rst
@@
respectively, where $g=2s+1$ is the spin multiplicity of the gas particle, and $h$ is the Planck constant. In addition, the characteristic size of the system can be determined by the fiducial length in numerical implementation,
@@
where $m_{\text{Pl}}=(\hbar c/G)^{1/2}$ is the Planck mass, $\lambda_{C}=\hbar/mc$ is the reduced Compton wavelength of the particle ($\hbar=h/2\pi$), and $\hat{r}$ is the dimensionless radius. Furthermore, we can express the dimensional quantities of the system using their corresponding dimensionless counterpart denoted with a “hat,” i.e., $n=(c^{2}/Gm\zeta^{2})\hat{n}$, $\rho=(c^{2}/G\zeta^{2})\hat{\rho}$, $u=(c^{4}/G\zeta^{2})\hat{u}$, $p=(c^{4}/G\zeta^{2})\hat{p}$, and $M=(c^{2}\zeta/G)\hat{M}$.
Figure 1: Normalized classical Maxwell–Boltzmann distribution (top) and quantum Fermi–Dirac distribution with $\alpha(R)=0$ (bottom). The cutoff energy is fixed at $bw=0.3$, and the boundary temperature varies as $b=k_{B}T(R)/mc^{2}=10^{-3}$ (blue), $10^{-2}$ (orange), and $10^{-1}$ (red). In the classical limit, the energy peak shifts toward higher energies as the temperature increases, whereas in the quantum limit, the trend reverses as the fermions become fully degenerate.
@@ III Fermi Degeneracy Pressure
In Fig. 1, we compare the normalized distribution function in the classical (top) and quantum (bottom) limits in the relativistic regime $bw=0.3$ for the boundary temperature $b=10^{-3}$ (blue), $10^{-2}$ (orange), and $10^{-1}$ (red). In the classical Maxwell-Boltzmann limit, most of the particles are non-relativistic with $bx\lesssim 10^{-2}$ for $b=10^{-3}$. As $b$ increases to $0.1$, more and more particles occupy the high energy state. Thus to trigger the instability in this case, one must increase the boundary temperature $b$ such that the majority of the particles can exist in the high energy state $bx\sim 0.1$. On the other hand, for the quantum limit $\alpha(R)=0$, the distribution is peaked at $bx\approx 0.3$ even at the low temperature $b=10^{-3}$. As $b$ increases further, the distribution changes mildly as in this limit the Pauli blocking effect dominates.
@@
Furthermore, we evaluate the velocity dispersion $v/c=\sqrt{3p/\rho c^{2}}$ and adiabatic index $\gamma=1+p/u$, where pressure $p$, energy density $\rho c^{2}$, and internal energy $u$ are given in Eq. 9. We consider the Maxwell–Boltzmann distribution and the Fermi–Dirac distribution with $\alpha(R)=\mu(R)/k_{B}T(R)=-10$ and 0, ranging from the classical to quantum limits. The $b$ values are $b=k_{B}T(R)/mc^{2}=10^{-3}$, $10^{-2}$, and $10^{-1}$, the same as those in Fig. 1.
@@ IV Dynamical Instability
In Fig. 3, we show the stable and unstable regions in the plane of the boundary temperature $b$ and the central cutoff $bw(0)$ for the degeneracy $\alpha(R)=-50$, 10-10, 5-5, and 0, covering the classical to quantum limits. In the classical limit ($\alpha(R)=-50$), the instability can only occur when the boundary temperature is high $b=k_{B}T(R)/mc^{2}>9\times 10^{-2}$ and the central cutoff is $bw(0)\gtrsim 0.35$ (relativistic regime, see Eq. 11). As the degeneracy $\alpha(R)$ increases to 10-10, another unstable region appears for $b<1.5\times 10^{-2}$ due to Fermi pressure, while the requirement on the central cutoff remains almost the same. In this case, the unstable region with high boundary temperature is still preserved, where the thermal pressure is the dominant source of the instability. Interestingly, there is a gap in the boundary temperature between the two unstable regions. In the gap region, both thermal pressure and Fermi pressure are relevant, but overall they are not strong enough to trigger the instability. As $\alpha(R)$ further increases, the gap shrinks and vanishes for $\alpha(R)\gtrsim-5$. For $\alpha(R)=0$, both regions join smoothly and the instability condition becomes insensitive to the boundary temperature $b$, as long as the central cutoff $bw(0)\gtrsim 0.3$.
@@
$\lambda_{\rm dB}(r)$ $=\lambda_{\rm C}\left[\left(1+\frac{3k_{B}T(r)}{2mc^{2}}\right)^{2}-1\right]^{-1/2}$ (14)
$=\lambda_{\rm C}\left[\left(1+\frac{3b/2}{1-bw(r)}\right)^{2}-1\right]^{-1/2},$
where $\lambda_{\rm C}=\hbar/mc$ is the reduced Compton wavelength. Using Eq. 10, we have the following relation
@@ V Black Hole Mass
With our numerical results as tabulated in Appendix C, we can calculate the critical mass for marginally stable configurations satisfying $\left<\gamma\right>=\gamma_{\rm cr}$. Since the dynamical evolution after reaching the instability leads to horizon formation Saijo:2002qt ; Shibata:2002br ; Shapiro:2002kk , we assume here that a black hole with the critical mass will form following the onset of instability. From Eq. 10, the gravitational mass of the bound sphere is given by
@@
When the degeneracy is large, Fermi pressure is the dominant source to trigger the instability even at a relatively low boundary temperature. In this regime, the critical mass is independent of the boundary temperature. In contrast, for the classical Maxwell–Boltzmann distribution, where thermal pressure is the only source of the instability and the only way to trigger the instability is to increase the boundary temperature. Note that for $\alpha(R)=-50$, the system is in the classical limit and follows the Maxwell–Boltzmann distribution, there are no marginally stable configurations in the low-temperature regime, as indicated in Fig. 3. We have checked that these features are generic and insensitive to the specific truncation model in Eq. 1.
Figure 4: Critical mass $M$ as a function of particle mass $m$ for the degeneracy $\alpha(R)=-50$ (left), 10-10 (middle), and 0 (right), obtained by varying the boundary temperature $b=0.5$ (red), $10^{-1}$ (orange), and $10^{-2}$ (blue). At the low temperature $b=10^{-2}$, marginally stable configurations are not found for $\alpha(R)=-50$, where the classical Maxwell–Boltzmann distribution is recovered, while the critical mass converges for $\alpha(R)=-10$ and 0 as the core becomes completely degenerate. The shaded gray band in each panel indicates the black hole mass range of astrophysical interest. The lower bound, $1\,{\rm M_{\odot}}$, corresponds to stellar-mass black holes, while the upper bound, $10^{11}{\rm\,M_{\odot}}$, approximates the most massive black holes discovered.
Fig. 4 shows the critical mass as a function of particle mass for $\alpha(R)=-50$ (left), 10-10 (middle), and 0 (right) for $b=0.5$ (red), $0.1$ (orange), and $10^{-2}$ (blue). For a given $\alpha(R)$, the critical mass increases with the boundary temperature $b$. For $\alpha(R)=-50$, the instability cannot be triggered at a low temperature $b=10^{-2}$, but this case has the highest critical mass at the high-temperature regime among the three examples. Thus in the classical regime, black hole seeds tend to be more massive. In the low-temperature regime $b=10^{-2}$, the critical mass converges and becomes insensitive to $\alpha(R)$. In this completely degenerate regime, the critical mass reaches the minimum for a given particle mass.
@@
Now we consider scenarios where dark matter clumps collapse into black holes due to gravothermal evolution, such as in the self-interacting dark matter framework Balberg:2001qg ; Feng:2020kxv ; Feng:2021rst . For fermonic dark matter, there is a lower bound on the particle mass, i.e., the Tremaine-Gunn bound Tremaine:1979we , below which dark matter halos cannot form. Recent studies found that the lower bound is $m\gtrsim 0.4\,{\rm~keV}/c^{2}$ Boyarsky:2008ju , although it could be relaxed if many distinct species of fermions are allowed Davoudiasl:2020uig . Taking $m=0.4\,{\rm keV/c^{2}}$ in Eq. 18, we find that the critical mass is $3.9\times 10^{12}\,{\rm M_{\odot}}$. This would exceed the mass of the most massive supermassive black holes observed to date, $\sim 10^{11}{\rm\,M_{\odot}}$ Brockamp:2016aa ; Dullo:2017aa . For a thermally produced dark matter particle, the gravitational lensing constraint from JWST data sets a lower limit on its mass of $6.1\,{\rm keV}/c^{2}$ Keeley:2024brx , corresponding to a critical mass of $1.67\times 10^{10}\,{\rm M_{\odot}}$ from Eq. 18. Conversely, if $\sim 10^{7}\,{\rm M_{\odot}}$ black holes associated with the JWST little red dots Matthee:2023utn ; Harikane:2023aa ; Kocevski:2023aa ; Pacucci:2023oci ; Maiolino:2023bpi are formed through the collapse of degenerate fermionic cores, the particle mass is estimated as $m\sim 100{\rm\,keV/c^{2}}$ Arguelles:2023hab ; Arguelles:2023kqw .